OpenStudy (anonymous):

Two cars raced at a race track. The faster car traveled 20 mph faster than the slower car. In the time that the slower car traveled 165 miles, the faster car traveled 225 miles. If the speeds of the cars remained constant, how fast did the slower car travel during the race?

OpenStudy (anonymous):

so, Let x= the speed of the faster car y = the speed of the slower car x = y + 30 (faster car traveled 20 mph faster than the slower car) 165/y = 225/x (since the time = distance / speed) The solution of the system of equations is x = 112.5 mph y = 82.5 mph The speed of the faster car is 112.5 mph. The speed of the slower car is 52.5 mph